A harmonic oscillator is made by using a .60 kg frictionless block and an Ideal spring of unknown force constant. The oscillator is found to have a period of 150s and a max speed of 2 m/s. Find (a) the force constant of the spring and (b) the amplitude of the oscillation.
Answer #1

Solution)

Given, m=0.60 kg

T=150 secs

For, SHM

We know, T=2*pi*sqrt(m/k)

So, k=4*pie^2*m/T^2

K=4*3.14^2*0.60/150^2=0.00105 N/m

=====

B) K.E=displacement

0.5 mv^2=0.5 kA^2

Sqrt(mv^2/k)=A^2

Sqrt(0.60*2^2/0.00105)=A

A=47.80 m (Ans)

=====